OpenStudy (angel_kitty12):

Can someone check my answer please? Thanks Two functions are shown f(x)=3x+1 g(x)=-3/2x-7 what is the value of x when f(x)=g(x)? for x, I got -4. Please check.

OpenStudy (angel_kitty12):

And if i did it wrong, could someone explain it to me? Thanks

OpenStudy (angel_kitty12):

Anyone?

OpenStudy (gabbyalicorn):

Does it have to be a whole number?

OpenStudy (gabbyalicorn):

http://prnt.sc/awwj5b What I got

OpenStudy (angel_kitty12):

I don't believe so

OpenStudy (angel_kitty12):

Thank you

OpenStudy (shubhipapa):

3x+1=-3/2x-7 3x+3/2x=-8 9/2x=-8 x=-8*2/9 x=-1.77